OpenStudy (anonymous):

Solve: 3x^2 + 16x - 9 = 3

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

I'll get you started 3x^2 + 16x - 9 = 3 3x^2 + 16x - 9 - 3 = 0 3x^2 + 16x - 12 = 0 Now use the quadratic formula to solve for x

OpenStudy (anonymous):

oh right! thank you @jim_thompson5910

jimthompson5910 (jim_thompson5910):

tell me what you get so I can check your answer

OpenStudy (anonymous):

ok im doing it now

OpenStudy (anonymous):

just to verify the equation its -16x +/- the square root of (16)^2-4(3)(-12)/2(3) correct?

OpenStudy (anonymous):

Solving 3x2+16x-12 = 0 by the Quadratic Formula According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 3 B = 16 C = -12 Accordingly, B2 - 4AC = 256 - (-144) = 400 Applying the quadratic formula : -16 ± √ 400 x = —————— 6 Can √ 400 be simplified ? The prime factorization of 400 is 2•2•2•2•5•5 To be able to remove something from under the radical, there have to be 2 instances of it because we are taking a square i.e. 2nd root. √ 400 = √ 2•2•2•2•5•5 =2•2•5•√ 1 = ± 20 • √ 1 = ± 20 So now we are looking at: x = ( -16 ± 20) / 6 Two real solutions: x =(-16+√400)/6 = u solve or: x =(-16-√400)/6 = u solve then i'll check to see if u get the concept :)

OpenStudy (anonymous):

i got -16-\[\sqrt{-2928}\] all over 6 but idk if that is the final answer

OpenStudy (anonymous):

x =(-16-√400)/6 = (-8-10)/3 = x =(-16+√400)/6 = (-8+10)/3 = Let me simplify, see if u can get it!

OpenStudy (anonymous):

@xokatexo do you need further help in understanding the concept?

OpenStudy (anonymous):

is that the final answer?

OpenStudy (mathmale):

The quadratic formula should "work" fine here, but I've found it easier to "factor by grouping." \[3x ^{2}+16x-2=3x ^{2}=18x-2x-2=3x(x+6)-2(x+6).\]

OpenStudy (anonymous):

No, u have to simplify that to find the final two solutions. I mean maybe i'm not explaining this right, sorry I'm better at solving then explaining. Are there any terms that can't be simplified ? no so u can simplify :)

OpenStudy (anonymous):

so my answer wasnt right

OpenStudy (anonymous):

where did you get 400 from @Nanalew

OpenStudy (mathmale):

(x+6) is the common factor here. Factoring it out, we get (x+6)(3x-2)=0. What are the solutions of that?

OpenStudy (anonymous):

Two real solutions: x =(-16+√400)/6 = (-8+10)/3 = 0.66667 or: x =(-16-√400)/6 = (-8-10)/3 = -6.00000 Here maybe understanding the answer will help u understand the problem. :D

OpenStudy (anonymous):

@mathmale i think im supposed to use the quadratic formula

OpenStudy (anonymous):

@Nanalew so this is what i did

OpenStudy (mathmale):

OK, so be it. Good luck!

OpenStudy (anonymous):

@agent0smith can u help me explain, i don't think I'm getting this concept across plus ur really good at explaining

OpenStudy (mathmale):

Be certain to use the quadratic formula on 3x^2+16x-12=0.

OpenStudy (anonymous):

3x^2+16x-9=3 -3 -3 --------------- 3x^2 +16x-12=0 -b+/- the square root of b^2-4ac -------------------------------- 2a -16 +/- the square root of (16)^2-4(3)(-12) -------------------------------------- 6

OpenStudy (mathmale):

Let @xokatexo be the judge of that.

OpenStudy (anonymous):

then i got -16+/- the square root of -2928 ---------------------------- 6 as my answer

OpenStudy (mathmale):

Sorry, but something's 'way off here. Whoever got 400 under the radical of the quadratic formula was correct. \[x=\frac{ -16+\sqrt{400} }{ 6 }\] is one solution. What is the other?

OpenStudy (anonymous):

okay i see where your coming from, here maybe @mathmale is on the right track it is easier to solve this way.. Simplify 3x2+16x-9 - 3 Trying to factor by splitting the middle term Factoring 3x2+16x-12 The first term is, 3x2 its coefficient is 3 The middle term is, +16x its coefficient is 16 The last term, "the constant", is -12 Step-1 : Multiply the coefficient of the first term by the constant 3 • -12 = -36 Step-2 : Find two factors of -36 whose sum equals the coefficient of the middle term, which is 16 -36 + 1 = -35 -18 + 2 = -16 -12 + 3 = -9 -9 + 4 = -5 -6 + 6 = 0 -4 + 9 = 5 -3 + 12 = 9 -2 + 18 = 16 Step-3 : Rewrite the polynomial splitting the middle term using the two factors found in step 2 above, (-2) and (18) 3x2 - 2x + 18x - 12 Step-4 : Add up the first 2 terms, pulling out like factors : x • (3x-2) Add up the last 2 terms, pulling out common factors : 6 • (3x-2) Now add up the four terms of step 3 : (x+6) • (3x-2) Solve 2.2 Solve : 3x-2 = 0 Add 2 to both sides of the equation : 3x = 2 Divide both sides of the equation by 3: x = 2/3 Solving a Single Variable Equation : 2.3 Solve : x+6 = 0 Subtract 6 from both sides of the equation : x = -6 See two solutions, is this easier to understand. It can be solved multiple ways :)

OpenStudy (mathmale):

Remember: "order of operations" is extremely important. Under the radical sign you have the following expression:\[16^{2}-4(3)(-12).\]

OpenStudy (mathmale):

According to "order of operations," we MUST multiply and/or divide BEFORE we add or subtract. Therefore, evaluate 4(3)(-12) BEFORE subtracting the result from 16^2. Try it.

OpenStudy (mathmale):

@xokatexo : I hope you're still with us!

OpenStudy (anonymous):

yeah i am im trying to do it that way

OpenStudy (anonymous):

Do my methods not make sense??

OpenStudy (anonymous):

its 256-144

OpenStudy (mathmale):

So, the radicand (the number under the square root operator) is .... ???? @xokatexo , you're on the right track except for a sign error. Mind going back and checking?

OpenStudy (anonymous):

OHHH okay i got it

OpenStudy (anonymous):

omg thanks so much wow ididnt pay attention to the other negative to make it positive

OpenStudy (anonymous):

ohhh when solving the quadratic formula here re-look at this step!! Accordingly, B2 - 4AC = 256 - (-144) = 400 now look back at the problem :)

OpenStudy (anonymous):

so its -16 + the square root of 400 ------------------------- 6

OpenStudy (mathmale):

I'm so glad y ou've both gotten the correct result! Now, would you please find the two solutions? x1=? x2=?

OpenStudy (anonymous):

YESSS ! :) let me repost :) Solving 3x2+16x-12 = 0 by the Quadratic Formula According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A In our case, A = 3 B = 16 C = -12 Accordingly, B2 - 4AC = 256 - (-144) = 400 Applying the quadratic formula : -16 ± √ 400 x = —————— 6 Can √ 400 be simplified ? The prime factorization of 400 is 2•2•2•2•5•5 To be able to remove something from under the radical, there have to be 2 instances of it because we are taking a square i.e. 2nd root. √ 400 = √ 2•2•2•2•5•5 =2•2•5•√ 1 = ± 20 • √ 1 = ± 20 So now we are looking at: x = ( -16 ± 20) / 6 Two real solutions: x =(-16+√400)/6 = (-8+10)/3 = 0.66667 or: x =(-16-√400)/6 = (-8-10)/3 = -6.00000 see now complete problem, do u understand ?

OpenStudy (anonymous):

\[-16+\sqrt{400}\] ---------------- 6

OpenStudy (anonymous):

and then -

OpenStudy (anonymous):

actually thank you so much you both helped me so much!!!!!!

OpenStudy (mathmale):

My great pleasure. Come back soon!

OpenStudy (anonymous):

medal?

OpenStudy (mathmale):

Awwww, Nanalew! We help others just through the goodness of our hearts, don't we? :)

OpenStudy (anonymous):

ugh i cant give you both medals but i became fans!!

OpenStudy (anonymous):

thanks @mathmale u helped me explain, here i medal u too. plus i worked really hard.... did u see my posts..... tell me u don't enjoy a reward for helping jeesh

OpenStudy (mathmale):

I'm just ribbing you. I'm not telling you not to enjoy a well-earned reward! :)

OpenStudy (anonymous):

yeah u seriously helped @mathmale, if u hadn't pointed out @xokatexo hadn't use order of operations i would have gone on babbling haha and yes i get it now, i'm still struggling w/ english so forgive me for not understanding humor

OpenStudy (anonymous):

happy to help and good luck to everyone :)

OpenStudy (agent0smith):

@xokatexo make sure you follow order of operations, because this is colossally wrong and will cause you a lot of trouble in algebra For evaluating: (16)^2-4(3)(-12) you did: 16^2 =256 256-4=252 252(-36)=-9072 http://www.mathsisfun.com/operation-order-pemdas.html

OpenStudy (anonymous):

yeah thanks!! idk i usually do im just blanking today haha

OpenStudy (anonymous):

Thanks @agent0smith u r the best at explaining ! :D

OpenStudy (anonymous):

thanks though!!!!